# Systematic analysis of the D-wave charmonium states with the QCD sum rules

### v3

@@ 1 Introduction
Charmonium is a flavorless meson made of a charmed quark and its antiquark, which is an ideal system for studying QCD up to the non-perturbative regime, and offers us an excellent topic to investigate strong interactions, therefore, systematically exploring the charmonium states can deepen our understanding of the dynamics involving the heavy quarknioum systems. Since the observation of the $J/\psi$ in 1974 [1, 2], experimental and theoretical physicists have been wrestling with the charmonium states. Prior to discovery of the $X(3872)$ in 2003 [3], many charmonium states, such as the $\eta_{c}$, $\eta_{c}^{\prime}$, $J/\psi$, $\psi^{\prime}$, $h_{c}$, $\chi_{c0}$, $\chi_{c1}$, $\chi_{c2}$, $\psi(3770)$, $\psi(4040)$, $\psi(4415)$, etc, were reported by the experimental collaborations [4, 5, 6, 7, 8, 9, 10], which can be classified in the traditional quark model. In fact, as early as the 1980s, the theoretical physicists began to focus on those particles and examined the mass spectrum, decay modes and collision processes in details [11, 12, 13, 14].
@@ 3 Numerical results and discussions
In order to avoid parameter dependence to the utmost extent, we prefer the central value of the pole contributions $55\%$ for all the charmonium states $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$. To avoid contaminations from the higher resonances and continuum states, we adopt the constraint $\sqrt{s_{0}}=M+0.45\sim 0.65\,\rm{GeV}$ for the continuum threshold parameters, and try to adjust the continuum threshold parameters to reproduce the charmonium masses $M$ via trial and error, as the largest energy gap between the ground state and first radial excited state is $0.65\,\rm{GeV}$ for the charmonium states (i.e. $M_{\eta_{c}^{\prime}}-M_{\eta_{c}}=0.65\,\rm{GeV}$) [56]. When the threshold $s=4m_{c}^{2}$ is reached, the endpoint singularity arise from the factor $s=4m_{c}^{2}$ in the denominator of the QCD spectral density. Due to the minor contribution from the gluon condensate terms, we implement the routine replacements $s=4m_{c}^{2}\to s=4m_{c}^{2}+\Delta^{2}$ with $\Delta^{2}=m_{c}^{2}$ in numerical computations, then the corresponding Borel parameters and masses are obtained.
@@
Figure 3: The contributions of the vacuum condensates of dimension $n$ with variations of the Borel parameters $T^{2}$, where the $A$, $B$, $C$ and $D$ denote $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$ charmonium states, respectively.
@@
The comparison of our mass predictions with those from other theoretical approaches, as summarized in Table 5. While potential models generally predict masses in the range of 3.773.843.77-3.84 GeV, consistent with our results, the Bethe-Salpeter equation and some potential models yield significantly different values, particularly for the $\psi_{2}$ and $\eta_{c2}$ states. Further experimental data are essential to advance our understanding of the D-wave charmonium systems. In particular, the observation of the $\eta_{c2}$ state and more precise measurements of masses, widths, and decay channels for the charmonia are crucial to test theoretical predictions.
@@ 4 Conclusion
In this work, we perform a systematic analysis of the conventional D-wave charmonium states $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$ in the framework of the QCD sum rules by carrying out the operator product up to the condensates of dimension-6, including both the gluon condensate $\langle\frac{\alpha_{s}GG}{\pi}\rangle$ and the three-gluon condensate $\langle g_{s}^{3}GGG\rangle$. Notably, despite significant differences in the values of the three-gluon condensate between the two parameter sets I and II adopted, the resulting mass predictions remain remarkably stable and consistent. This robustness underscores the reliability of the QCD sum rules and the dominant role of perturbative contributions. We observe that the predicted masses favor identifying the $\psi(3770)$, $\psi_{2}(3823)$ and $\psi_{3}(3842)$ as the D-wave charmonium states $1{}^{3}D_{1}$, $1{}^{3}D_{2}$ and $1{}^{3}D_{3}$, respectively, as the $\eta_{c2}(1{}^{1}D_{2})$ has not been observed experimentally yet, we make prediction to be confronted to experimental data in the future. Meanwhile, we can take the decay constants as elementary input parameters to study the strong decays and radiative decays to make the identifications in a more robust way.

### v4

@@ 1 Introduction
Charmonium, a flavorless meson made of a charmed quark and its antiquark, provides an ideal system for studying QCD up to the non-perturbative regime. Its study offers us an excellent topic to investigate strong interactions. Therefore, systematically exploring the charmonium states can deepen our understanding of the dynamics involving the heavy quarknioum systems. Since the observation of the $J/\psi$ in 1974 [1, 2], experimental and theoretical physicists have actively wrestling with the charmonium states. Prior to the discovery of the $X(3872)$ in 2003 [3], many charmonium states, such as the $\eta_{c}$, $\eta_{c}^{\prime}$, $J/\psi$, $\psi^{\prime}$, $h_{c}$, $\chi_{c0}$, $\chi_{c1}$, $\chi_{c2}$, $\psi(3770)$, $\psi(4040)$, $\psi(4415)$, etc, were reported by the experimental collaborations [4, 5, 6, 7, 8, 9, 10], which can be classified in the traditional quark model. In fact, as early as the 1980s, the theoretical physicists began to focus on those particles and examined the mass spectrum, decay modes and collision processes in detail [11, 12, 13, 14].
@@ 3 Numerical results and discussions
In order to avoid parameter dependence to the utmost extent, we prefer the central value of the pole contributions $55\%$ for all the charmonium states $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$. To avoid contaminations from the higher resonances and continuum states, we adopt the constraint $\sqrt{s_{0}}=M+0.45\sim 0.65\,\rm{GeV}$ for the continuum threshold parameters, and try to adjust the continuum threshold parameters to reproduce the charmonium masses $M$ via trial and error, as the largest energy gap between the ground state and first radial excited state is $0.65\,\rm{GeV}$ for the charmonium states (i.e. $M_{\eta_{c}^{\prime}}-M_{\eta_{c}}=0.65\,\rm{GeV}$) [58]. The selection of the continuous threshold parameters both refer to the experimental results, and require ensuring that the operator product expansion at QCD side, the hadron sides and QCD sides are matched through dispersion relation to get the properties of the hadron states. The primary criteria are internal stability of the QCD sum rules, pole dominance, convergence of the OPE, and the existence of the Borel platform. These criteria take precedence over importing precise energy gaps predicted by other models, because the QCD sum rules extract hadron properties directly from correlation functions without assuming a specific dynamical model.
In the QCD spectral densities given in Eqs.(9)-(12), when the threshold $s=4m_{c}^{2}$ is reached, the endpoint singularity arise from the factor the QCD spectral densities appear in the gluon condensate $\frac{1}{\sqrt{s(s-4m_{c}^{2})}}$. Due to the minor contribution from the gluon condensate terms compared to the perturbative term as shown in Fig.3, we attempt to implement the routine replacements $s=4m_{c}^{2}\to s=4m_{c}^{2}+\Delta^{2}$ with $\Delta^{2}=\frac{m_{c}^{2}}{2}$, $m_{c}^{2}$, $\frac{3m_{c}^{2}}{2}$ in numerical computations. The determination of numerical results for masses at three different $\Delta^{2}$ values are consistent. To ensure as much consistency in parameters selection as possible, we adopted the $\Delta^{2}=m_{c}^{2}$ substitution as used in our previous work [18, 19], then the corresponding Borel parameters and masses are obtained.
@@
Figure 3: The contributions of the vacuum condensates of dimension $n$ with variations of the Borel parameters $T^{2}$ of the input parameters are set to central values, where the $A$, $B$, $C$ and $D$ denote $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$ charmonium states, respectively.
@@
uncertainty $\Delta x_{i}$ $\Delta M_{\psi_{1}(1{}^{3}D_{1})}$ $\Delta M_{\psi_{2}(1{}^{3}D_{2})}$ $\Delta M_{\eta_{c2}(1{}^{1}D_{2})}$ $\Delta M_{\psi_{3}(1{}^{3}D_{3})}$
$m_{c}$ $\pm 0.025\rm{GeV}$ $\pm 22.07\rm{MeV}$ $\pm 23.31\rm{MeV}$ $\pm 22.58\rm{MeV}$ $\pm 22.24\rm{MeV}$
$\sqrt{s_{0}}$ $\pm 0.1\rm{GeV}$ $\pm 54.54\rm{MeV}$ $\pm 52.55\rm{MeV}$ $\pm 53.04\rm{MeV}$ $\pm 53.75\rm{MeV}$
$T^{2}$ $\pm 0.3\rm{GeV}$ $\pm 26.95\rm{MeV}$ $\pm 29.11\rm{MeV}$ $\pm 28.85\rm{MeV}$ $\pm 24.99\rm{MeV}$
Table 5: The mass uncertainties of the parameters variation for input parameter I.
uncertainty $\Delta x_{i}$ $\Delta M_{\psi_{1}(1{}^{3}D_{1})}$ $\Delta M_{\psi_{2}(1{}^{3}D_{2})}$ $\Delta M_{\eta_{c2}(1{}^{1}D_{2})}$ $\Delta M_{\psi_{3}(1{}^{3}D_{3})}$
$m_{c}$ $\pm 0.025\rm{GeV}$ $\pm 21.92\rm{MeV}$ $\pm 23.51\rm{MeV}$ $\pm 22.69\rm{MeV}$ $\pm 22.15\rm{MeV}$
$\sqrt{s_{0}}$ $\pm 0.1\rm{GeV}$ $\pm 54.36\rm{MeV}$ $\pm 52.45\rm{MeV}$ $\pm 53.06\rm{MeV}$ $\pm 53.62\rm{MeV}$
$T^{2}$ $\pm 0.3\rm{GeV}$ $\pm 26.78\rm{MeV}$ $\pm 26.67\rm{MeV}$ $\pm 29.30\rm{MeV}$ $\pm 23.95\rm{MeV}$
$\langle\frac{\alpha_{s}GG}{\pi}\rangle$ $\pm 0.004\rm{GeV}^{4}$ $\pm 0.22\rm{MeV}$ $\pm 1.31\rm{MeV}$ $\pm 0.61\rm{MeV}$ $\pm 2.61\rm{MeV}$
$\langle g_{s}^{3}GGG\rangle$ $\pm 0.385\rm{GeV}^{6}$ $\pm 1.45\rm{MeV}$ $\pm 0.33\rm{MeV}$ $\pm 1.49\rm{MeV}$ $\pm 1.68\rm{MeV}$
Table 6: The mass uncertainties of the parameters variation for input parameter II.
Our prediction $M_{\psi_{1}}=3.77\pm{0.09}\,\rm{GeV}$ favors assigning the conventional $\psi(3770)$ (rather than the $\mathcal{R}(3780)$ reported by the BESIII collaboration [16, 17]) as the D-wave charmonium $\psi_{1}$. Our result of $M_{\psi_{1}(1{}^{3}D_{1})}$ supports the dominance of the D-wave Fock component for $\psi(3770)$. The mass of $\psi(3770)$ lies significantly above $\psi(2S)$. Although the consideration of S-D mixing effects on vector charmonia can interpret their experimental di-electric widths [61], the S-wave admixture were very large, such agreement would be unlikely. The predicted mass $M_{\psi_{2}}=3.82\pm{0.09}\,\rm{GeV}$ is consistent with the observation of the Belle, BESIII and LHCb collaborations [20, 21, 22, 23, 24], and favors the identification $\psi_{2}$ indeed. While the $1{}^{1}D_{2}$ charmonium state has not been discovered yet, we estimate the mass to be $M_{\eta_{c2}}=3.83\pm{0.09}\,\rm{GeV}$, which is expected to be observed at the BESIII, LHCb, Belle II collaborations in the future. Furthermore, we obtain the mass $M_{\psi_{3}}=3.84\pm{0.08}\,\rm{GeV}$, and tentatively identify the $X(3842)(\psi_{3}(3842))$ as the $1{}^{3}D_{3}$ charmonium state [25]. We can take the decay constants $f_{\psi_{1}/\psi_{2}/\eta_{c2}/\psi_{3}}$ are elementary input parameters to study the strong decays and radiative decays to make the identifications in a more robust way.
The comparison of our mass predictions with those from other theoretical approaches, as summarized in Table 7. The potential models generally predict masses in the range of 3.773.843.77-3.84 GeV, if we consider the uncertainty ranges, the predictions from these methods fall within our calculated error bands. Therefore, the following discussion will focus on comparison of the central values, for the $\psi_{1}$ state, the quark model predictions are generally higher than our results, which may be attributed to inherent model dependencies. Our central values of the unobserved $\eta_{c2}$ is closed to some quark models. This concordance between the methods is particularly noteworthy and reinforces the credibility of approaches. Further experimental data are essential to advance our understanding of the D-wave charmonium systems. In particular, the observation of the $\eta_{c2}$ state and more precise measurements of masses, widths, and decay channels for the charmonia are crucial to test theoretical predictions.
@@
The decay constant is a crucial parameter in the calculation of strong, leptonic, semileptonic, and radiative decays. For instance, in Ref.[62], the decay constants $f_{\psi_{1}}$, $f_{\psi_{2}}$, $f_{\eta_{c2}}$, $f_{\psi_{3}}$ presented in this work is employed within the framework of three-point correlation functions to compute the semileptonic decay width of D-wave charmonia. There is a scarcity of research dedicated to direct computate the decay constants for D-wave charmonia [63]. Our calculated constants are not merely numerical outputs but are key phenomenological parameters that bridge the gap between mass predictions and the future analysis of decay widths and branching ratios.
The values of the non-perturbative vacuum condensates, particularly the triple-gluon condensate $\langle g_{s}^{3}GGG\rangle$, are not as precisely determined as the fundamental parameters like the charm quark mass $m_{c}$. Different extraction methods from QCD sum rules analyse or phenomenological fits can lead to differing numerical estimates. The predicted masses for all four D-wave charmonium states remain remarkably stable and consistent between the two parameter sets, as detailed in Tables 3 and 4, the variations are well within the quoted uncertainties. This demonstrates that our mass predictions are dominated by the perturbative contribution and are not sensitive to the specific choice of the gluon condensate values and the triple-gluon condensate. While the masses are robust, we observe the minor sensitivity in the decay constants, the parameter sets I and II are considered.
@@ 4 Conclusion
In this work, we perform a systematic analysis of the conventional D-wave charmonium states $\psi_{1}$, $\psi_{2}$, $\eta_{c2}$ and $\psi_{3}$ in the framework of the QCD sum rules. We carry out the operator product up to the condensates of dimension-6, which includes both the gluon condensate $\langle\frac{\alpha_{s}GG}{\pi}\rangle$ and the three-gluon condensate $\langle g_{s}^{3}GGG\rangle$. Notably, despite significant differences between the values of the three-gluon condensate in the two adopted parameter sets I and II adopted, the resulting mass predictions remain remarkably stable and consistent. This robustness underscores the reliability of the QCD sum rules and the dominant role of the perturbative contributions. We observe that the predicted masses favor identifying the $\psi(3770)$, $\psi_{2}(3823)$ and $\psi_{3}(3842)$ as the D-wave charmonium states $1{}^{3}D_{1}$, $1{}^{3}D_{2}$ and $1{}^{3}D_{3}$, respectively. Since the $\eta_{c2}(1{}^{1}D_{2})$ has not been observed experimentally yet, we make prediction to be confronted to experimental data in the future. Meanwhile,the obtained decay constants can serve as elementary input parameters in subsequent study of the strong decays, leptonic and radiative decays to make the identifications in a more robust way.